OpenStudy (anonymous):

How many grams of magnesium metal will react completely with 5.2 liters of 4.0 M HCl? Show all of the work needed to solve this problem. Mg (s) + 2 HCl (aq) --> MgCl2 (aq) + H2 (g)

OpenStudy (abb0t):

Start with the information of what you're given. Do you know the units of M? I believe it is mol/L? Yes

OpenStudy (abb0t):

Use dimensional analysis to cancel out the units of Liters so that you're left with units of mol! Do you follow me?

OpenStudy (anonymous):

Yeah, I know about molarity, but starting it off is my problem.

OpenStudy (abb0t):

This process is very similar to how you did in algebra: \(\sf \color{}{\frac{a}{2bc} \times \frac{c}{2}=\frac{a}{4b}}\) right? Because the two \(c\)'s cancel out.

OpenStudy (anonymous):

So this would be the framework?

OpenStudy (abb0t):

Yessir. Cancel out the get the units mol. Then, you can go from mol HCl \(\rightarrow\) mol Mg\(^{2+}\) and if you notice, you have a 2:1 ratio, so you're going to be dividing by two.

OpenStudy (anonymous):

But where would the parts go, necessarily?

OpenStudy (anonymous):

I did m = M(v) and got 20.8 moles

OpenStudy (abb0t):

\(\sf \color{red}{5.2~L \times \frac{4.0 ~mol~HCl}{L} \times \frac{1~mol~Mg}{2~mol~HCl}\times molar ~mass~of~ Mg^{2+}}\)

OpenStudy (anonymous):

Got it. So 5.21* 4.0 /2 = 10.42?

OpenStudy (abb0t):

Now multiply it by the molar mass of magnesium to get grams! You can do that by looking at your periodic table. Remember that they are in units of grams per mol. So make sure you know what units your caanceling out. And you're done

OpenStudy (anonymous):

Would it look like 10.42 mol Mg * (24.31g Mg/ 1 mol) = 235.31???

OpenStudy (abb0t):

Yes sir.

OpenStudy (anonymous):

Thank you very much!